# Capturing Turbulence with Numerical Dissipation: a Simple Dynamical Model for Unresolved Turbulence in Hydrodynamic Simulations

### v2

@@ 3.3.1 Mean small-scale turbulent energy
Figure 4: Comparison of coarse-grained DNS (dashed lines) with the SLES resimulations at corresponding resolutions (colored lines). The top panel shows the evolution of the global Mach number, $\mathcal{M}=\sqrt{2E_{\rm kin}}$ in the adopted units ($\langle\rho\rangle\equiv 1$). The bottom panel shows the mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, normalized by the initial total kinetic energy in the box, $E_{\rm kin,0}$. The SLES model captures both the evolution of $E_{\rm turb}$ and its scale dependence. As shown in Appendix A, $E_{\rm turb}$ is underestimated for 1283 due to lack of convergence on these small scales; however, we still show the results for this grid to demonstrate that the scaling of $E_{\rm turb}$ with the grid size remains the same in SLES.
As pointed out in the previous section, the fact that the SLES model reproduces the temporal evolution of the mean small-scale turbulent energy, $E_{\rm turb}$, in the DNS is a non-trivial test of the model, as the timesteps spaced by more than a box-crossing time can be viewed as independent realizations. In decaying turbulence, the global Mach number of turbulence decreases with time, and therefore different times also probe different Mach-number regimes. As the top panel of Figure 4 shows, the Mach number decreases from $\sim 8.5$ to $\lesssim 2$ after 8 eddy crossing times of the simulation box, $t_{\rm cross,0}$. The SLES model performs well in these different regimes.
@@
The shape of $E_{\rm turb}$ evolution is qualitatively similar on different scales and is only offset to higher values for larger scales (i.e., smaller $N^{3}$). This is because larger scales contain more kinetic energy. The SLES model captures this dependence on scale. In this model, the information about scale is contained both in the implicit source term—as the numerical dissipation rate depends on the resolution—and in the explicit turbulence decay term which is proportional to the local cell-crossing time of subgrid turbulence (Equation (3)).
One noticeable deviation of the SLES resimulations from the DNS is the slightly higher global Mach number and mean $E_{\rm turb}$ at $t\sim 1\text{--}3\times t_{\rm cross,0}$. As was shown in the previous section, such a feature can be suppressed by modifying the model for the subgrid turbulence decay (see Figure 3). However, this feature might also be due to the initial transient associated with the smoothing of initial sharp cell-to-cell density and velocity gradients resulting from the coarse-graining high-resolution ICs. We checked that this feature does not depend on how the subgrid $e_{\rm turb}$ is initialized. In these tests we initialize as $e_{\rm turb}=0$ throughout the box; however, if we start from the $e_{\rm turb}$ equal to the coarse-grained ICs values using Equation (8), this feature still persists. Such a transient can be avoided by testing the model in a continuously driven turbulence box setup, which we leave to a separate study.
@@ 3.3.2 Turbulent velocity–density correlation
Figure 6: Top panel: Comparison of the $\rho$$\sigma$ relation between the coarse-grained DNS (dashed black lines) and low-resolution SLES resimulation (solid blue lines). Bottom set of panels: dependence of scatter in $\sigma$ on gas compression rate (negative divergence of the resolved velocity field, $-\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$) at three fixed densities indicated with green lines in the top panel and labeled in the corner of each panel. Each of the panels shows cells with densities bound by 1.5 times higher and 1.5 times lower than these values. For reference, the dotted red lines show the medians from the low-resolution resimulations with an explicit LES model proposed by schmidt11. The SLES model qualitatively reproduces the trend of $\sigma$ with $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$ although it does show a steeper increase at high compression rates. This behavior, however, is analogous to that in a more complex explicit LES model.
@@
As Figure 2 showed, local turbulent energy on small scales is correlated with the gas density. Figure 5 demonstrated this correlation quantitatively for both the DNS coarse-grained using a $32^{3}$ grid (left column) and an SLES resimulation at the corresponding resolution (right column). The rows correspond to different times, from top to bottom: 1, 3, and $5$ global turbulence crossing time of the simulation box, $t_{\rm cross,0}$. The range of the $y$-axes changes from top to bottom to compensate for the decrease in the normalization of this relation due to turbulence decay; to help visualize this decreasing trend, the red line is kept the same in all panels (see below).
@@
First, focusing on the DNS, the $\rho$$\sigma$ relation exhibits several remarkable features both in slope and scatter. The slope is shallower at earlier times and at higher $\sigma$, implying that turbulent energy changes isothermally with density. At later times and at lower $\sigma$, the slope steepens and becomes close to the adiabatic value, $\sigma\propto\rho^{1/3}$ (i.e., $e_{\rm turb}=\rho\sigma^{2}/2\propto\rho^{5/3}$), shown with the red thin line. This change is the most noticeable in the bottom panel where the median $\sigma_{\rm turb}$ exhibits a change in slope at $\rho\sim 0.5$. Overall, the changes in the slope are subtle indicating that, on average, small-scale turbulence remains in the intermediate regime.
@@
As the color shows, the scatter correlates with the divergence of resolved velocity, i.e., gas compression rate. This reflects the adiabatic heating (cooling) of turbulence as contracting (expanding) gas exerts the $PdV$ work on subgrid turbulence (robertson12). The magnitude of the compression rate decreases with time as the turbulence decays but the correlation of scatter with $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$ persists.
@@
A comparison of the right and left columns in the figure shows that the SLES model captures the slope, the magnitude of scatter, and its dependence on compression rate remarkably well. Such an agreement is particularly remarkable given how simple the model is: it involves only one tunable parameter that was fixed effectively by matching the normalization of this relation in one of the snapshots (see Section 3.2). The slope, the scatter, and their evolution are the predictions of the model, and they match the DNS results.
@@
To further quantify the agreement between the SLES model and DNS, the $\rho$$\sigma$ relations from both simulations are over-plotted in the top panel of Figure 6. The bottom set of panels shows the dependence of $\sigma$ on the resolved compression rate at three fixed values of $\rho=0.3$, 1, and 3. The SLES model captures the overall increase of $\sigma$ at higher compression rates, $-\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$, although it produces a steeper relation at high values of $-\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$. However, given the simplicity of the model, the agreement is still remarkable. For reference, the thin dotted line shows the results of a low-resolution simulation with an explicit LES model adopting the non-linear schmidt11 closure for the supersonic turbulence source (see Section 2.2). The performance of the SLES model is comparable to that of the explicit LES model without requiring assumptions about the closure relation for the turbulence source term.
@@ 3.3.3 Power spectra
As the dashed line in the top panel shows, in the DNS, the velocity power spectrum is a power law with a constant slope of $\approx-5/3$ down to $k/k_{\rm min}\sim 100$ and falls off at higher $k$. Given the resolution of the DNS, $1024^{3}$, the fall-off at these scales indicates that, in the ART code, numerical dissipation becomes important on the scales of $\sim 10$ grid cells. A similar numerical dissipation scale was found by kritsuk11 for analogous second-order hydrodynamic solvers (e.g., the RAMSES code; ramses). It is worth noting that the power spectrum compensated by the $k^{-5/3}$ slope exhibits a local maximum (i.e., the bottleneck effect) at $k/k_{\rm min}\sim 40$ indicating that numerical dissipation effects kick in on $\sim 1024/40\sim 25$ cells scale (kitsionas09; federrath10; federrath11), which can also explain the excess of power at low $k$ in the SLES runs.
@@
As the middle panel of Figure 7 shows, the corresponding fall-off also appears in the density power spectrum. This fall-off also shifts to larger scales (smaller $k$) with decreasing resolution. The density power spectrum falls off with $k$ roughly as $P(\rho,k)\propto k^{-1.1}$. This slope is within the range reported by federrath13 for the non-magnetized turbulence runs with Mach numbers $\lesssim 10$ before the gas self-gravity is turned on (see also federrath10).
@@
Apart from having qualitatively similar slopes, the drop-off of the $\rho$ and $e_{\rm turb}$ power spectra at high $k$ exhibit similar dependences on the resolution. This is another manifestation of the $\rho$$\sigma$ correlation (see Section 3.3.2). Indeed, due to this correlation, $e_{\rm turb}=\rho\sigma^{2}/2$ can be viewed as a passive marker of density and therefore it will follow the dependence of $\rho$ on resolution.
@@
To summarize, despite its simplicity, the SLES model captures many non-trivial features of small-scale turbulence in a direct simulation. It can correctly predict the global behavior of the mean turbulent energy, its dependence on scale, as well as the local correlations between the turbulent energy, gas density, and compression rate. In the next section, we investigate the model performance in a more practical application, a simulation of a disk galaxy.
@@ 4 Galaxy formation simulations
As an example of practical application, in this section, we explore the performance of the SLES model in a galaxy formation simulation. The model is used to dynamically predict the distribution of locally variable star-formation efficiency that depends on the properties of turbulence on unresolved scales. As we will show, the simple SLES model produces results close to a more complex explicit LES model.
@@ 4.1 Galaxy model and methods
Figure 8: Face-on view of the same galaxy re-simulated with our new semi-implicit LES (top row) and explicit LES (bottom panel) models for unresolved turbulence. Columns show, from left to right: slices of gas density ($n\equiv\rho/m_{\rm p}$; for simplicity, we assume that the average molecular weight equals one), subgrid turbulent velocity ($\sigma_{\rm turb}\equiv\sqrt{2e_{\rm turb}/\rho}$), locally variable star formation efficiency per freefall time ($\epsilon_{\rm ff}$; see Equation 10), and the surface density of SFR computed using star particles younger than 100 Myr. To guide the eye, thin lines indicate density isocontours corresponding to $n=20\;{\rm cm^{-3}}$. Overall, both models produce similar distributions of $\sigma_{\rm turb}$, $\epsilon_{\rm ff}$, and resulting $\dot{\Sigma}_{\rm\star}$. The most noticeable differences are the higher levels of turbulence on the disk outskirts and in the center in the former, which originate partially due to the filtering of the bulk shear in the explicit model.
We initialize an isolated $L_{\star}$ disk galaxy simulation using one of the snapshots from the fiducial run of semenov17; semenov18; semenov19, which was started from the initial conditions from the AGORA code comparison project (agora-intro; agora-disk). The computational mesh is adaptively refined by splitting the cells when their gas mass exceeds $\sim 8300\;M_{\odot}$ or their local Jeans length (including the contributions from thermal and turbulent pressure) is resolved by fewer than 16 cells until the minimal cell size of $\Delta=40\;{\rm pc}$ is reached. Gas cooling and heating are modeled using the gnedin12 model, assuming constant solar metallicity and a uniform radiation background corresponding to the H2 photodissociation rate in the Lyman–Werner bands of $10^{-10}\;{\rm s^{-1}}$ (stecher67), which is suppressed in the dense gas following a shielding model calibrated against radiative transfer simulations of the ISM (the “L1a” model from safranekshrader17). Stellar feedback is implemented by injecting density-, metallicity-, and resolution-dependent radial momentum and thermal energy calibrated by martizzi15, with the momentum boosted by a factor of 5 to account for the effects of SN clustering (e.g., gentry17; gentry19) and cosmic rays (diesing18), assuming chabrier03 IMF.
In this paper, we rerun these simulations adopting different models for unresolved turbulence. The unresolved turbulence is dynamically tracked to predict the local dynamical state of gas and, coupled with it, locally variable star formation efficiency. Specifically, we use the common parametrization of the local SFR density via the star formation efficiency per freefall time,
@@
The virial parameter accounts for both the thermal and subgrid turbulent pressure as $\sigma_{\rm tot}\equiv\sqrt{c_{\rm s}^{2}+\sigma_{\rm turb}^{2}}$, with $\sigma_{\rm turb}\equiv\sqrt{2e_{\rm turb}/\rho}$ is modeled using the subgrid turbulence model for $e_{\rm turb}$.
In this paper, for the comparison with the performance of the SLES model, we use our fiducial subgrid turbulence model with an explicit eddy-viscosity closure that we extensively tested in previous works (semenov16; semenov17; semenov18; semenov19; semenov21-tf), henceforth, the explicit LES model. The turbulence source term in this closure depends on the gradients of the fluctuating component of the velocity field with the bulk flow velocity filtered out, the so-called “shear-improved” closure (leveque07; schmidt14). This is done to avoid spurious production of turbulence in constant shearing flows like differentially rotating disks.
In contrast to the shear-improved LES closure, the implicit source term in our SLES model provided by local numerical dissipation intrinsically depends on the gradients of the total velocity field. Although the two models produce similar results (as presented below), the differences between the models can be partially explained by the shear being removed from our explicit LES model. Indeed, the resimulation of our LES run with the shear filtering turned off (i.e., with the closure relation dependent on the total velocity like in SLES) shows results intermediate between our semi-implicit and explicit models. Here we only present the comparison with the shear-improved model because this is the explicit model that is best suited for applications in galaxy simulations (see Section 5 for discussion).

### v3

@@ 3.3.1 Mean Small-scale Turbulent Energy
Figure 4: Comparison of coarse-grained DNS (dashed lines) with the SLES resimulations at corresponding resolutions (colored lines). The top panel shows the evolution of the global Mach number, $\mathcal{M}=\sqrt{2E_{\rm kin}}$ in the adopted units ($\langle\rho\rangle\equiv 1$). The bottom panel shows the mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, normalized by the initial total kinetic energy in the box, $E_{\rm kin,0}$. The SLES model captures both the evolution of $E_{\rm turb}$ and its scale dependence. As shown in Appendix A, $E_{\rm turb}$ is underestimated for 1283 due to a lack of convergence on these small scales; however, we still show the results for this grid to demonstrate that the scaling of $E_{\rm turb}$ with the grid size remains the same in SLES.
As pointed out in the previous section, the fact that the SLES model reproduces the temporal evolution of the mean small-scale turbulent energy, $E_{\rm turb}$, in the DNS is a nontrivial test of the model, as the time steps spaced by more than a box-crossing time can be viewed as independent realizations. In decaying turbulence, the global Mach number of turbulence decreases with time, and therefore, different times also probe different Mach number regimes. As the top panel of Figure 4 shows, the Mach number decreases from $\sim 8.5$ to $\lesssim 2$ after 8 eddy crossing times of the simulation box, $t_{\rm cross,0}$. The SLES model performs well in these different regimes.
@@
The shape of $E_{\rm turb}$ evolution is qualitatively similar on different scales and is only offset to higher values for larger scales (i.e., smaller $N^{3}$). This is because larger scales contain more kinetic energy. The SLES model captures this dependence on scale. In this model, the information about scale is contained both in the implicit source term—as the numerical dissipation rate depends on the resolution—and in the explicit turbulence decay term, which is proportional to the local cell-crossing time of subgrid turbulence (Equation (3)).
One noticeable deviation of the SLES resimulations from the DNS is the slightly higher global Mach number and mean $E_{\rm turb}$ at $t\sim 1\text{--}3\times t_{\rm cross,0}$. As was shown in the previous section, such a feature can be suppressed by modifying the model for the subgrid turbulence decay (see Figure 3). However, this feature might also be due to the initial transient associated with the smoothing of the initial sharp cell-to-cell density and velocity gradients resulting from the coarse-graining high-resolution ICs. We checked that this feature does not depend on how the subgrid $e_{\rm turb}$ is initialized. In these tests we initialize as $e_{\rm turb}=0$ throughout the box; however, if we start from the $e_{\rm turb}$ equal to the coarse-grained ICs values using Equation (8), this feature still persists. Such a transient can be avoided by testing the model in a continuously driven turbulence box setup, which we leave to a separate study.
@@ 3.3.2 Turbulent Velocity–Density Correlation
Figure 6: Top panel: comparison of the $\rho$$\sigma$ relation between the coarse-grained DNS (dashed black lines) and low-resolution SLES resimulation (solid blue lines). Bottom set of panels: dependence of scatter in $\sigma$ on gas compression rate (negative divergence of the resolved velocity field, $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$) at three fixed densities indicated with green lines in the top panel and labeled in the corner of each bottom panel. Each of the panels shows cells with densities bound by 1.5 times higher and 1.5 times lower than these values. For reference, the dotted red lines show the medians from the low-resolution resimulations with an explicit LES model proposed by schmidt11. The SLES model qualitatively reproduces the trend of $\sigma$ with $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$ although it does show a steeper increase at high compression rates. This behavior, however, is analogous to that in a more complex explicit LES model.
@@
As Figure 2 shows, local turbulent energy on small scales is correlated with the gas density. Figure 5 demonstrates this correlation quantitatively for both the DNS coarse-grained using a $32^{3}$ grid (left column) and an SLES resimulation at the corresponding resolution (right column). The rows correspond to different times: from top to bottom, 1, 3, and $5$ global turbulence crossing times of the simulation box, $t_{\rm cross,0}$. The range of the $y$-axis changes from top to bottom to compensate for the decrease in the normalization of this relation due to turbulence decay; to help visualize this decreasing trend, the red line is kept the same in all panels (see below).
@@
First, focusing on the DNS, the $\rho$$\sigma$ relation exhibits several remarkable features both in slope and in scatter. The slope is shallower at earlier times and at higher $\sigma$, implying that turbulent energy changes isothermally with density. At later times and at lower $\sigma$, the slope steepens and becomes close to the adiabatic value, $\sigma\propto\rho^{1/3}$ (i.e., $e_{\rm turb}=\rho\sigma^{2}/2\propto\rho^{5/3}$), shown with the red thin line. This change is the most noticeable in the bottom panel, where the median $\sigma_{\rm turb}$ exhibits a change in slope at $\rho\sim 0.5$. Overall, the changes in the slope are subtle, indicating that, on average, small-scale turbulence remains in the intermediate regime.
@@
As the color shows, the scatter correlates with the divergence of resolved velocity, i.e., the gas compression rate. This reflects the adiabatic heating (cooling) of turbulence as contracting (expanding) gas exerts the $PdV$ work on subgrid turbulence (robertson12). The magnitude of the compression rate decreases with time as the turbulence decays, but the correlation of scatter with $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$ persists.
@@
A comparison of the right and left columns in the figure shows that the SLES model captures the slope, the magnitude of scatter, and its dependence on compression rate remarkably well. Such agreement is particularly remarkable given how simple the model is: it involves only one tunable parameter that is fixed effectively by matching the normalization of this relation in one of the snapshots (see Section 3.2). The slope, the scatter, and their evolution are the predictions of the model, and they match the DNS results.
@@
To further quantify the agreement between the SLES model and DNS, the $\rho$$\sigma$ relations from both simulations are overplotted in the top panel of Figure 6. The bottom set of panels shows the dependence of $\sigma$ on the resolved compression rate at three fixed values of $\rho=0.3$, 1, and 3. The SLES model captures the overall increase of $\sigma$ at higher compression rates, $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$, although it produces a steeper relation at high values of $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$. However, given the simplicity of the model, the agreement is still remarkable. For reference, the thin dotted line shows the results of a low-resolution simulation with an explicit LES model adopting the nonlinear schmidt11 closure for the supersonic turbulence source (see Section 2.2). The performance of the SLES model is comparable to that of the explicit LES model without requiring assumptions about the closure relation for the turbulence source term.
@@ 3.3.3 Power Spectra
As the dashed line in the top panel shows, in the DNS, the velocity power spectrum is a power law with a constant slope of $\approx-5/3$ down to $k/k_{\rm min}\sim 100$ and falls off at higher $k$. Given the resolution of the DNS, $1024^{3}$, the falloff at these scales indicates that, in the ART code, numerical dissipation becomes important on the scales of $\sim 10$ grid cells. A similar numerical dissipation scale was found by kritsuk11 for analogous second-order hydrodynamic solvers (e.g., the RAMSES code; ramses). It is worth noting that the power spectrum compensated by the $k^{-5/3}$ slope exhibits a local maximum (i.e., the bottleneck effect) at $k/k_{\rm min}\sim 40$ indicating that numerical dissipation effects kick in on $\sim 1024/40\sim 25$ cell scales (kitsionas09; federrath10; federrath11), which can also explain the excess of power at low $k$ in the SLES runs.
@@
As the middle panel of Figure 7 shows, the corresponding falloff also appears in the density power spectrum. This falloff also shifts to larger scales (smaller $k$) with decreasing resolution. The density power spectrum falls off with $k$ roughly as $P(\rho,k)\propto k^{-1.1}$. This slope is within the range reported by federrath13 for the nonmagnetized turbulence runs with Mach numbers $\lesssim 10$ before the gas self-gravity is turned on (see also federrath10).
@@
Apart from having qualitatively similar slopes, the drop-off of the $\rho$ and $e_{\rm turb}$ power spectra at high $k$ exhibits similar dependences on the resolution. This is another manifestation of the $\rho$$\sigma$ correlation (see Section 3.3.2). Indeed, due to this correlation, $e_{\rm turb}=\rho\sigma^{2}/2$ can be viewed as a passive marker of density, and therefore, will follow the dependence of $\rho$ on resolution.
@@
To summarize, despite its simplicity, the SLES model captures many nontrivial features of small-scale turbulence in a direct simulation. It can correctly predict the global behavior of the mean turbulent energy, its dependence on scale, and the local correlations between the turbulent energy, gas density, and compression rate. In the next section, we investigate the model performance in a more practical application, a simulation of a disk galaxy.
@@ 4 Galaxy Formation Simulations
As an example of practical application, in this section, we explore the performance of the SLES model in a galaxy formation simulation. The model is used to dynamically predict the distribution of locally variable star-formation efficiency that depends on the properties of turbulence on unresolved scales. As we will show, the simple SLES model produces results close to those of a more complex explicit LES model.
@@ 4.1 Galaxy Model and Methods
Figure 8: Face-on view of the same galaxy resimulated with our new SLES (top row) and explicit LES (bottom panel) models for unresolved turbulence. Columns show, from left to right, slices of gas density ($n\equiv\rho/m_{\rm p}$; for simplicity, we assume that the average molecular weight equals 1), subgrid turbulent velocity ($\sigma_{\rm turb}\equiv\sqrt{2e_{\rm turb}/\rho}$), locally variable star formation efficiency per freefall time ($\epsilon_{\rm ff}$; see Equation 10), and the surface density of SFR computed using star particles younger than 100 Myr. To guide the eye, thin lines indicate density isocontours corresponding to $n=20\;{\rm cm^{-3}}$. Overall, both models produce similar distributions of $\sigma_{\rm turb}$, $\epsilon_{\rm ff}$, and resulting $\dot{\Sigma}_{\rm\star}$. The most noticeable differences are the higher levels of turbulence on the disk outskirts and in the center in the former, which originate partially due to the filtering of the bulk shear in the explicit model.
We initialize an isolated $L_{\star}$ disk galaxy simulation using one of the snapshots from the fiducial run of semenov17; semenov18; semenov19, which was started from the initial conditions from the AGORA code comparison project (agora-intro; agora-disk). The computational mesh is adaptively refined by splitting the cells when their gas mass exceeds $\sim 8300\;M_{\odot}$ or their local Jeans length (including the contributions from thermal and turbulent pressure) is resolved by fewer than 16 cells until the minimal cell size of $\Delta=40\;{\rm pc}$ is reached. Gas cooling and heating are modeled using the gnedin12 model, assuming constant solar metallicity and a uniform radiation background corresponding to the H2 photodissociation rate in the Lyman–Werner bands of $10^{-10}\;{\rm s^{-1}}$ (stecher67), which is suppressed in the dense gas following a shielding model calibrated against radiative transfer simulations of the ISM (the “L1a” model from safranekshrader17). Stellar feedback is implemented by injecting density-, metallicity-, and resolution-dependent radial momentum and thermal energy calibrated by martizzi15, with the momentum boosted by a factor of 5 to account for the effects of SN clustering (e.g., gentry17; gentry19) and cosmic rays (diesing18), assuming the chabrier03 initial mass function.
In this paper, we rerun these simulations adopting different models for unresolved turbulence. The unresolved turbulence is dynamically tracked to predict the local dynamical state of gas and, coupled with it, locally variable star formation efficiency. Specifically, we use the common parameterization of the local star formation rate (SFR) density via the star formation efficiency per freefall time,
@@
The virial parameter accounts for both the thermal and subgrid turbulent pressure as $\sigma_{\rm tot}\equiv\sqrt{c_{\rm s}^{2}+\sigma_{\rm turb}^{2}}$, with $\sigma_{\rm turb}\equiv\sqrt{2e_{\rm turb}/\rho}$ modeled using the subgrid turbulence model for $e_{\rm turb}$.
In this paper, for comparison with the performance of the SLES model, we use our fiducial subgrid turbulence model with an explicit eddy-viscosity closure that we extensively tested in previous works (semenov16; semenov17; semenov18; semenov19; semenov21-tf), henceforth, the explicit LES model. The turbulence source term in this closure depends on the gradients of the fluctuating component of the velocity field with the bulk flow velocity filtered out, the so-called “shear-improved” closure (leveque07; schmidt14). This is done to avoid spurious production of turbulence in constant shearing flows like differentially rotating disks.
In contrast to the shear-improved LES closure, the implicit source term in our SLES model provided by local numerical dissipation intrinsically depends on the gradients of the total velocity field. Although the two models produce similar results (as presented below), the differences between the models can be partially explained by the shear being removed from our explicit LES model. Indeed, the resimulation of our LES run with the shear filtering turned off (i.e., with the closure relation dependent on the total velocity as in SLES) shows results intermediate between our semi-implicit and explicit models. Here we only present the comparison with the shear-improved model because this is the explicit model that is best suited for applications in galaxy simulations (see Section 5 for discussion).